Jordan Boat Company makes inexpensive aluminum fishing boats. Production is seasonal, with considerable activity occurring in the spring and summer. Sales and production tend to decline in the fall and winter months. During 2019, the high point in activity occurred in June when it produced 206 boats at a total cost of $149,870. The low point in production occurred in January when it produced 39 boats at a total cost of $48,000. Required Use the high-low method to estimate the amount of fixed cost incurred each month by Jordan Boat Company. Determine the total estimated cost if 160 boats are made.
Answer #1

Variable cost under high low method = (Cost at highest production activity-Cost at lowest production activity) / (Highest activity-Lowest activity).

= (149,870-48,000)/(206-39)

= 101,870/167

= 610 per boat

Total cost = Fixed + Variable

149,870 = Fixed + (610*206)

Fixed cost = 24,210

Estimated cost for 160 boats = (160*610)+24,210

= 121,810